OpenStudy (anonymous):

Graph the function and identify the vertical asymptote y=(1/x-4)-2

OpenStudy (anonymous):

if this is the function your referring to: \(\large y=\frac{1}{x-4}-2 \) then this graph looks exactly the same as: \(\large y=\frac{1}{x} \) with the following transformations: the graph of \(\large y=\frac{1}{x-4}-2 \) is a horizontal translation 4 units to the right and a vertical translation 2 units down.